OpenStudy (anonymous):

The lengths of the sides of a 6-sided polygon are an arithmetic sequence. The length of the shortest side is 3 meters. If the length of the next longer side is 5 meters, what is the length of the longest side?

OpenStudy (tkhunny):

Arithmetic Sequence - Difference between any two consecutive values is the same. 3 3 + 2 = 5 5 + 2 = 7 7 + 2 = 9 One more to go.

OpenStudy (tkhunny):

Okay, maybe two more to go.